OpenStudy (anonymous):

What would be all the seventh roots of unity? Help would be great! <3

OpenStudy (kc_kennylau):

\[\text{1. }\cos0+i\sin0=1\]\[\text{2. }\cos\left(\frac{360}7\right)+i\sin\left(\frac{360}7\right)\]\[\text{3. }\cos\left(2\times\frac{360}7\right)+i\sin\left(2\times\frac{360}7\right)\]\[\text{4. }\cos\left(3\times\frac{360}7\right)+i\sin\left(3\times\frac{360}7\right)\]\[\text{5. }\cos\left(4\times\frac{360}7\right)+i\sin\left(4\times\frac{360}7\right)\]\[\text{6. }\cos\left(5\times\frac{360}7\right)+i\sin\left(5\times\frac{360}7\right)\]\[\text{7. }\cos\left(6\times\frac{360}7\right)+i\sin\left(6\times\frac{360}7\right)\]

OpenStudy (anonymous):

Okay.. I see my answers were a bit different.. but thanks for your help! (:

OpenStudy (kc_kennylau):

what're your answers?